### v1

# Krait: A Backdoor Attack Against Graph Prompt Tuning

Abstract
Graph prompt tuning has emerged as a promising paradigm to effectively transfer general graph knowledge from pre-trained models to various downstream tasks, particularly in few-shot contexts. However, its susceptibility to backdoor attacks, where adversaries insert triggers to manipulate outcomes, raises a critical concern. We conduct the first study to investigate such vulnerability, revealing that backdoors can disguise benign graph prompts, thus evading detection. We introduce Krait, a novel graph prompt backdoor. Specifically, we propose a simple yet effective model-agnostic metric called label non-uniformity homophily to select poisoned candidates, significantly reducing computational complexity. To accommodate diverse attack scenarios and advanced attack types, we design three customizable trigger generation methods to craft prompts as triggers. We propose a novel centroid similarity-based loss function to optimize prompt tuning for attack effectiveness and stealthiness. Experiments on four real-world graphs demonstrate that Krait can efficiently embed triggers to merely 0.15% to 2% of training nodes, achieving high attack success rates without sacrificing clean accuracy. Notably, in one-to-one and all-to-one attacks, Krait can achieve 100% attack success rates by poisoning as few as 2 and 22 nodes, respectively. Our experiments further show that Krait remains potent across different transfer cases, attack types, and graph neural network backbones. Additionally, Krait can be successfully extended to the black-box setting, posing more severe threats. Finally, we analyze why Krait can evade both classical and state-of-the-art defenses, and provide practical insights for detecting and mitigating this class of attacks.
I Introduction
Graph neural networks (GNNs) have achieved exceptional success in various real-world applications, such as recommendation systems [recomm_app], anomaly detection [anomaly_detect], drug discovery, and disease diagnosis [bio_health]. Nevertheless, traditional GNNs heavily rely on manual annotations and are prone to over-fitting, especially when testing graphs are out-of-distribution. To address these issues, the “Graph pre-training and fine-tuning” paradigm has been widely employed. It pre-trains GNNs on diverse graph datasets and then transfers the learned graph knowledge to downstream tasks with minimal adaptation steps. Despite its success, fine-tuning large-scale GNNs can be time-consuming and resource-intensive. Additionally, pre-training tasks can misalign with downstream tasks in terms of semantic knowledge, learning objectives, and task difficulty, leading to negative transfers [all_in_one, graph_prompt_survey]. Recently, the “Graph pre-training and prompt-tuning” [GPPT, graph_prompts, all_in_one] has emerged as a promising solution to mitigate these misalignments. It utilizes graph prompts as free parameters to reformulate downstream tasks into pre-training tasks, thereby reducing computational overhead and adaptation complexity.
Graph Backdoor. Previous research has shown that GNNs are vulnerable to graph backdoors [graph_backdoor, unnoticeable_graph_backdoor, multi_gnn_backdoor], which implant triggers into graphs so that GNNs misclassify nodes attached with triggers as a target label while maintaining the accuracy of clean nodes. However, existing trigger generation methods are often random or computationally intensive [random_backdoor, graph_backdoor, explain_backdoor], reducing the efficacy of backdoor attacks. Additionally, Dai et al. [unnoticeable_graph_backdoor] empirically demonstrate that most graph backdoors consume large attack budgets, i.e., poisoning rate (PR), to guarantee attack effectiveness on large-scale graphs, but their generated triggers can be easily detected by homophily and node-similarity-based defense mechanisms. Besides, existing studies primarily focus on one-to-one attacks, with more complex all-to-one and all-to-all attacks less explored. Although a few efforts have been made to design one-to-many and many-to-one attacks [multi_gnn_backdoor, one_to_many_backdoor], the word “many” in these contexts only refers to a few labels. Most importantly, to the best of our knowledge, backdoor attacks against graph prompt tuning are yet to be explored. However, prior work cannot be trivially adapted into graph prompt tuning since the majority of them are designed for traditional GNNs [unnoticeable_graph_backdoor, random_backdoor, explain_backdoor, multi_gnn_backdoor]. Even though several studies support backdooring pre-trained GNNs, poisoning the pre-training graphs can be time-consuming, and perturbing pre-trained GNNs [graph_backdoor, pre-train_backdoor, contrastive_backdoor] will inherit the limitations of the pre-trained models themselves. Due to privacy protection and proprietary concerns, many companies and institutions cannot open-source their pre-training datasets along with the pre-training process, making it infeasible to poison the pre-training phase of graph prompt tuning.
Graph Prompt Backdoor. We observe an interesting phenomenon: graph prompt tuning and graph backdoor attacks both append small-scale subgraphs to the input graphs, but serve for inverse purposes. Graph prompt tuning treats subgraphs as graph prompts to improve the performance of downstream tasks, while graph backdoor attacks implant subgraphs as triggers to selected poisoned candidates, ensuring they are classified as the desired label [graph_backdoor, unnoticeable_graph_backdoor]. This inherent similarity motivates us to investigate the following three questions: 1) Are graph prompt tuning models susceptible to backdoor attacks? 2) If so, can attackers craft graph prompts as triggers and exploit this vulnerability to facilitate graph backdoor attacks? 3) How can we design an effective and stealthy graph prompt backdoor for various attack types? Notably, designing such a graph prompt backdoor poses additional challenges. For one thing, graph prompt tuning typically adapts the pre-trained GNNs to downstream tasks within a few epochs, which might not establish a solid mapping between the triggers and target labels, thus compromising the effectiveness of backdoor attacks. For another, disguising triggers as benign graph prompts requires attackers to generate adaptive triggers, otherwise, these triggers can be destroyed during the prompt tuning process.
Practicability. Recently, an open-source library called ProG [^1] https://github.com/sheldonresearch/ProG/tree/main has been released to the public, which packs many state-of-the-art (SOTA) graph prompt tuning models [graph_prompt_survey]. Due to its flexibility, it is anticipated that more developers can tailor this library to the specific requirements of their Graph-Prompt-Tuning-as-a-Service (GPTaaS) APIs. However, developers can also become malicious GPTaaS providers, who may poison the customized downstream graph with well-crafted triggers before model training, or modify handcrafted or auto-generated graph prompts to implement graph backdoors. Additionally, malicious users of GPTaaS APIs can manipulate the handcrafted graph prompts and inject them through user interfaces, or even send malicious instructions to generate graph prompts for poisoned samples. Furthermore, there is a promising trend of developing large language model (LLM)-powered graph foundation models [graph_llm]. Malicious users can interact with LLM interfaces and manipulate graph prompts to poison graph foundation models. Vulnerabilities in LLMs could also be exploited through malicious prompts [prompt_as_trigger1, prompt_as_trigger2], further compromising the functionality of graph foundation models. Moreover, multiple prompts have been utilized to adapt LLMs for specific tasks, such as recommendations [multi_prompt_recommendation], or to enhance their accuracy and robustness for general tasks [multi_prompt]. This approach inherently increases the risk of concealing triggers within one of the multiple graph prompts, offering a more stealthy attack channel for graph prompt backdoors.
Our Proposal. In this paper, we conduct the first study to examine the vulnerability of graph prompt tuning against backdoor attacks. We demonstrate that graph prompt tuning is susceptible to these security threats and propose two node-level homophily metrics and three centroid similarity-based metrics to investigate graph prompt tuning before and after backdoor injections. Our observations reveal consistent patterns between graph prompt tuning and graph backdoors, highlighting the inherent vulnerability to crafting graph prompts as triggers. To tackle the aforementioned challenges, we introduce Krait (Backdoor against graph prompt tuning), a novel graph prompt backdoor. Inspired by the “venom optimization” theory [venom], we design a model-agnostic metric, label non-uniformity homophily, to identify the most vulnerable poisoned candidates before model training, which significantly reduces computational costs. Guided by the Chinese idiom “using somebody’s spear to attack his shield”, we exploit the inherent vulnerability in graph prompt tuning to disguise triggers as benign prompts. These triggers are then adaptively tuned along with benign graph prompts, thus reducing the risk of detection and destruction. Furthermore, we strictly set PRs between $0.15\%$ and $2\%$, making Krait more evasive to existing defense mechanisms. We develop three trigger generation methods tailored to various attack scenarios and advanced attack types under the more stringent white-box setting. Additionally, we incorporate a novel centroid similarity-based loss to improve the effectiveness and stealthiness of Krait. Finally, we extend Krait to the black-box setting, enhancing its practicality.
Contributions. Our five key contributions are as follows:
  • We propose two node-level homophily metrics and three centroid similarity-based metrics to analyze the behaviors of graph prompt tuning before and after backdoor injections. The observed patterns unveil the intrinsic vulnerability to treating graph prompts as triggers.
  • We introduce Krait, a novel graph prompt backdoor attack comprising three key components: 1) a model-agnostic met- ric called label non-uniformity homophily to identify the most vulnerable nodes as poisoned candidates; 2) three versatile trigger generation methods to generate graph prompts as triggers; 3) a novel centroid similarity-based loss function to boost the attack’s effectiveness and stealthiness.
  • Our extensive experiments show that Krait can generate effective and covert triggers across various transfer cases, attack types, trigger generation methods, and GNN backbones under the more stringent white-box setting. Remarkably, Krait only poisons 0.15%-2% of training nodes for different attack types. In some specific cases, Krait only implants minimal-scale triggers into $2$ nodes but achieves a 100% attack success rate (ASR).
  • We successfully extend Krait to the black-box setting.
  • We demonstrate that existing defenses cannot effectively detect and mitigate Krait, highlighting the urgent need for more reliable defenders against graph prompt backdoors. We further provide general suggestions for researchers and practitioners to safeguard graph prompt tuning models.
Social Impact. Our research underscores the significant security threats posed by graph prompt backdoors, particularly in the context of LLM-powered graph foundation models that have become prevalent in diverse real-world applications. In high-stake scenarios such as healthcare [healthcare_graph_prompt], attackers could mislead GPTaaS APIs to classify patients who should be diagnosed with illnesses as non-disease cases, potentially delaying critical treatments and endangering lives. Similarly, in less sensitive areas such as recommendation systems [prompt_recommend], attackers could manipulate the systems to expose more inappropriate or harmful content, leading to severe social impacts such as exacerbating discrimination, fostering hatred, and aiding criminal activities. We bring these critical security concerns into the research community, hoping that our work will stimulate future research on detecting and mitigating graph prompt backdoors. We will release our code to support these efforts.
II Background
II-A Notations
Given an undirected attributed graph $\mathcal{G}=(\mathcal{V},\mathcal{E},X)$, $\mathcal{V}$ denotes a node set with $|\mathcal{V}|$ nodes and each node $v$ is associated with a feature vector $X_{v}\in\mathcal{R}^{1\times d}$, where $d$ is the dimension of node features, $\mathcal{E}$ represents an edge set with $|\mathcal{E}|$ edges, which can be captured by the adjacency matrix $A\in\mathcal{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ and $A_{u,v}=1$ i.i.f. the edge $(u,v)\in\mathcal{E}$. Hereby, we reform $\mathcal{G}=(A,X)$.
II-B Graph Neural Networks
In essence, the remarkable performance of the majority of GNNs root in message-passing mechanisms, which aggregate each node $v\in\mathcal{V}$’s information from its local neighborhood $\mathcal{N}(v)$ and further update its node embedding $H^{l}_{v}$ after $l$ iterations. Formally, this process can be expressed as:
$H_{v}^{l}=UPD^{l}(H_{v}^{l-1},AGG^{l-1}(\{H_{u}^{l-1}:u\in\mathcal{N}(v)\}))$ (1)
where $H_{v}^{0}=X_{v}$, and $UPD$ and $AGG$ are two arbitrary differentiable functions to design diverse GNN variants. For node classification tasks, generally, $H_{v}^{l}$ is fed into a linear classifier with a softmax function for final prediction. As for graph classification tasks, readout functions can be deployed to encode graph-level representation by $Z_{\mathcal{G}}=REA({H_{u}^{l}|u\in\mathcal{V}})$, and $REA$ can be a simple permutation invariant function, e.g., average, summation or a more sophisticated graph pooling operation. Since transferring graph-level knowledge is more effective and generalizable, we reformulate node-level tasks into graph-level tasks by constructing $k$-hop ego-networks $\mathcal{G}_{s(v)}$ for $v\in\mathcal{V}$ [all_in_one], where the $k$-hop ego-network denotes the induced subgraph of the target node $v$ and the shortest path distance from any other node to $v$ is set to $k$.
II-C Graph Pre-training and prompt tuning
Pre-training Phase. Graph pre-training transfers general graph knowledge to downstream tasks without requiring manual annotations. As indicated in [graph_prompt_survey], existing graph pre-training methods can be categorized into two main branches: graph generative learning and graph contrastive learning (GCL)-based methods. We solely consider the latter throughout this paper. Generally, GCL-based methods exploit graph augmentation techniques to generate two augmented views $\hat{\mathcal{G}}_{s(v),1}$ and $\hat{\mathcal{G}}_{s(v),2}$ for each $\mathcal{G}_{s(v)}$ as positive pairs, where $\hat{\mathcal{G}}_{s(v)}\sim q(\hat{\mathcal{G}}_{s(v)}|\mathcal{G}_{s(v)})$, and $q(\cdot|\mathcal{G}_{s}(v))$ denotes the augmentation distribution conditioned on $\mathcal{G}_{s(v)}$, $v\in\mathcal{V}$; negative pairs $\hat{\mathcal{G}}_{{s(v)}^{\prime}}$ are generated from the rest $|\mathcal{V}|-1$ augmented graphs [graphcl]. GCL-based graph pre-training aims to minimize the contrastive loss to pre-train the GNN model $f_{pre}$:
$\min\limits_{\theta_{pre}}\mathcal{L}_{pre}=\sum\limits_{v\in\mathcal{V}}log\frac{-exp(sim(Z_{\mathcal{G}_{s(v),1}},Z_{\mathcal{G}_{s(v),2}})/\tau)}{\sum_{v\prime=1,v\prime\neq v}exp(sim(Z_{\mathcal{G}_{s(v),1}},Z_{\mathcal{G}{s(v)\prime,2}})/\tau)}$ (2)
where $\tau$ is the temperature parameter.
Prompt-tuning Phase. Graph prompts generally consist of three components: prompt tokens, token structures, and inserting patterns [graph_prompt_survey]. The simplest graph prompts are additional feature vectors concatenated to existing node features, while more advanced designs consider the inner structures of prompt tokens and their integration with the input graphs [all_in_one]. Graph prompt tuning flexibly utilizes such graph prompts to reformulate downstream tasks into the pre-training task space [graph_prompt_survey, all_in_one]. As for prompt generation, the graph prompt tuning model $f_{prm}$ employs a graph prompt generator $g_{\psi}:\mathcal{G}\rightarrow\mathcal{G}$ to transform each ego-network $\mathcal{G}_{s(v)}$ into the prompted subgraph $g_{\psi}(\mathcal{G}_{s(v)})=\mathcal{G}_{s(v)}\oplus\mathcal{G}_{\Delta}$, where $\mathcal{G}_{\Delta}$ are learnable graph prompts with prompt tokens $\mathcal{P}$ and token structures $\mathcal{E}_{\Delta}$, and $\oplus$ denotes inserting patterns. As for task-specific tuning, given a frozen pre-trained GNN $f_{pre}$, $f_{prm}$ strives to minimize the classification loss between final prediction $\hat{Y}$ and ground-truth label $Y$, which can be formulated as:
$\min\limits_{\psi,\theta}\mathcal{L}_{prm}=\sum\limits_{v\in\mathcal{V}}l(f_{prm}(g_{\psi}(\mathcal{G}_{s(v)}),\theta_{nor}),Y_{i})$ (3)
where $\theta_{nor}$ contains parameters from the graph prompt tuning model $\theta_{prm}$ and frozen parameters from the pre-trained model $\theta_{pre}$. $l(\cdot)$ can be any loss function, e.g. entropy loss.
II-D Graph Backdoor
Different from existing graph backdoors [graph_backdoor] that inject triggers to the pre-trained GNN model $f_{pre}$, our attackers endeavor to generate a universal trigger $\delta$ and embed it to the graph prompt tuning model $f_{prm}$. They aim to misclassify the target nodes attached with triggers as the desired target label, and meanwhile, maintain a high classification accuracy of clean samples. which can be expressed as:
$\begin{split}\min\limits_{\theta_{tri}}\mathcal{L}_{tri}&=\sum\limits_{v\in\mathcal{V}_{c}}l(f_{prm}(g_{\psi}(\mathcal{G}_{s(v)}),\theta),Y_{v})+\\ &\phantom{=\;\;}\sum\limits_{u\in\mathcal{V}_{p}}l(f_{prm}(g_{\psi}(\mathcal{G}_{s(u)})\oplus\delta,\theta),Y_{t})\end{split}$ (4)
where $Y_{t}$ is the desired target label and $\theta$ combines $\theta_{nor}$ and $\theta_{tri}$. $\mathcal{V}_{c}$ and $\mathcal{V}_{p}$ denote clean and poisoned node sets.
We consider a more stringent white-box setting where attackers are only allowed to modify customized downstream graphs, e.g., feature modification, edge addition/removal, and label flipping; without access to original pre-training graphs and the corresponding pre-training process. Attackers can obtain intermediate training results and attach triggers to the testing nodes during the inference phase. They can interact with graph prompts through a specific user interface or by gaining partial control of the downstream task training. We further extend our attack into the black-box setting. More details will be presented in Section LABEL:sec:framework.
III Motivation and Problem Statement
In this section, we design two types of metrics—node-centric homophily [unnoticeability_node_centric_homo] and centroid similarity—to investigate the behaviors of graph prompt tuning before and after backdoor injections. Our experimental results demonstrate that graph prompt tuning is vulnerable to backdoor attacks and exposes a new attack channel for disguising triggers as benign graph prompts. These findings motivate us to develop more effective and stealthy backdoor attacks against graph prompt tuning. Finally, we formalize our research problem in a new subsection.
We use All-in-One, the state-of-the-art (SOTA) graph prompt tuning model [all_in_one], to conduct our experiments. All-in-One supports the customization of downstream graphs, making it more flexible and practical for real-world applications. Additionally, it provides a universal framework for graph prompt tuning, thus enhancing its generalizability. Notably, many other graph prompt tuning models [GPPT, VNT] can be viewed as special instances of All-in-One [graph_prompt_survey].
We consider two transfer settings, within-domain transfer and out-of-domain transfer. We employ four real-world datasets from diverse domains, namely, Cora, CiteSeer [cora_citeseer], Physics, and Computers [physics_computers]. We select the small-scale datasets, Cora and Citeseer, as the target domains, and all four graphs as the source domains. Please refer to Appendix LABEL:appendix_dataset for detailed graph description and statistics.
III-A Metric Design
In this subsection, we introduce two node-centric homophily metrics and three centroid similarity metrics to investigate the behaviors of graph prompt tuning before and after backdoor injections.
III-A1 Node-centric Homophily
Generally, homophily describes the phenomenon where connected nodes tend to share similar labels or features [unnoticeability_node_centric_homo]. Since many real graphs are homophilic and most existing GNNs implicitly follow homophily assumptions, we leverage homophily to investigate the behaviors of graph prompt tuning before and after backdoor injections.
However, graph prompt tuning generally constructs an ego-network for each targeted node, which involves transforming nodes into local subgraphs. Each ego-network is paired with a specific node label, but the nodes within the ego-network do not have labels, which is also the case with the graph prompt appended to each ego-network. Therefore, we cannot directly calculate label-based homophily. To address this issue, based on node-centric homophily [unnoticeability_node_centric_homo], we propose two metrics—local-subgraph homophily and global-view homophily—to compute node feature similarity between a node and its $1$-hop neighbors in different ways.
Definition 1.
(Local-subgraph Homophily). We first compute the node-centric homophily for each node $u\in\mathcal{G}_{s(v)}$ and then take the average to represent the subgraph-level homophily of $\mathcal{G}_{s(v)}$, i.e., the local homophily of each node $v\in\mathcal{G}$, which can be formally expressed as:
$h_{v}=\frac{1}{|\mathcal{V}_{s(v)}|}\sum\limits_{u\in\mathcal{G}_{s(v)}}sim(r_{u},X_{u}),r_{u}=\sum\limits_{w\in\mathcal{N}(u)}\frac{1}{\sqrt{d_{u}d_{w}}}X_{w}$ (5)
where $|\mathcal{V}_{s(v)}|$ is the node size of the ego-network $\mathcal{G}_{s(v)}$, $d_{u}$ is the degree of node $u$, and $sim(\cdot)$ is the predefined similarity metric, e.g., cosine similarity.